OpenStudy (anonymous):

Solve 7x − 2y = −3 14x + y = 14

OpenStudy (anonymous):

14x+y=14 y=14-14x y=14(1-x) put above value of y on another equation 7x-2(14(1-x))=-3 7x-28+28x=-3 35x=28-3 35x=25 x=5/7 putting x in any above equation to get value of y 7(5/7)-2y=-3 5-2y=-3 y=4